OpenStudy (anonymous):

WILL MEDAL AND FAN. please help!! Find the value of c that makes the expression a perfect square. write the expression as the square of a binomial 1. y^2-14y+c 2.s^2+17s+c 3. z^2+24z+c

OpenStudy (anonymous):

1. c=49 2. c=72.25 3. c=144 fan and medal??

OpenStudy (anonymous):

\[y^2-2*y*7+...^2=(y-..)^2\]

OpenStudy (here_to_help15):

@shehryar no direct answer can you provide @maggie148 with an explanation?

OpenStudy (here_to_help15):

\[answers* \]

OpenStudy (anonymous):

@shehryar yea i will if you can help exlpain.i kinda want to learn it. I have a final exam coming up. thank you!

OpenStudy (anonymous):

u have to keep in mind a simple formula (a+b)2=a2+2(a)(b)+b2

OpenStudy (here_to_help15):

@shehryar http://openstudy.com/code-of-conduct you might want to read this again

OpenStudy (anonymous):

i am sorry @Here_to_Help15 will keep in mind next time

OpenStudy (here_to_help15):

Thanks :)